[Brahms-l] Shift report 20050212 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sat Feb 12 2005 - 00:04:32 EST
1) No beam at begining of shift. Good Store from about 18:00 to 23:00
<br>2) Run No.   FS@3<sup>o</sup> MRS@40<sup>o</sup> 
<br>--------------------------------------------
<br>r13111     272k    106k
<br>r13112     315k    136k
<br>r13113     230k    115k 
<br>r13114     210k    130k 
<br>3)At the start of the shift, there was a problem turning on the <br>cathods on T3,T4,T5 from BRAHVO. Radik found that the detectors tripped and cleared the trips.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Feb 12 00:04:51 2005

This archive was generated by hypermail 2.1.8 : Sat Feb 12 2005 - 00:05:04 EST